OpenStudy (anonymous):

how do i solve 5n-2+3n=5

OpenStudy (anonymous):

Begin by adding the like terms, so add 5n and 3n. now you have 8n-2=5

OpenStudy (anonymous):

5n-2+3n=5 +2 +2 5n+3n=7 What is 5n+3n=???? 8n=7 now divide both sides by 8.